# Optimal Stopping for Sequential Bayesian Experimental Design

### v1

@@ Abstract
In sequential Bayesian experimental design, the number of experiments is usually fixed in advance. In practice, however, campaigns may terminate early, raising the fundamental question: when should one stop? Threshold-based rules are simple to implement but inherently myopic, as they trigger termination based on a fixed criterion while ignoring the expected future information gain that additional experiments might provide. We develop a principled Bayesian framework for optimal stopping in sequential experimental design, formulated as a Markov decision process where stopping and design policies are jointly optimized. We prove that the optimal rule is to stop precisely when the immediate terminal reward outweighs the expected continuation value. To learn such policies, we introduce a policy gradient method, but show that naïve joint optimization suffers from circular dependencies that destabilize training. We resolve this with a curriculum learning strategy that gradually transitions from forced continuation to adaptive stopping. Numerical studies on a linear-Gaussian benchmark and a contaminant source detection problem demonstrate that curriculum learning achieves stable convergence and outperforms vanilla methods, particularly in settings with strong sequential dependencies.
Keywords Bayesian experimental design $\cdot$ Sequential experimental design $\cdot$ Optimal stopping $\cdot$ Markov decision process $\cdot$ Policy gradient
@@ 1 Introduction
Sequential experimentation plays a central role in science and engineering, from clinical trials (Murphy, 2003) and materials discovery (Lookman et al., 2019) to environmental monitoring (Krause et al., 2008). In these settings, practitioners adaptively choose new experiments based on outcomes of earlier ones. A fundamental yet often overlooked question is: when should experimentation stop?
@@
The prevailing practice is to fix the number of experiments in advance. However, real-world campaigns frequently face uncertain budgets, limited resources, or diminishing returns that make early termination desirable. Simple threshold-based stopping rules are often used in practice, halting experimentation once a predefined criterion (e.g., posterior variance, remaining budget) falls below a cutoff. While convenient, such rules are inherently myopic: they depend only on the current state and ignore the expected long-term trade-off between the value of additional information and the cost of further experiments. As a result, thresholds can terminate too early, missing valuable insights, or too late, wasting resources (see Figure 1).
@@
Figure 1: Accumulated reward trajectories over six experiments, illustrating the limitations of threshold-based stopping. Here, a fixed threshold of 1.8 (dashed line) yields very different outcomes across trajectories: the orange path is stopped prematurely after one experiment, missing the higher rewards achievable at stages 2–3; the blue path never crosses the threshold, leading to unnecessary continuation past its peak; the green path happens to cross the threshold near its maximum, but only by chance. Threshold rules are thus highly outcome-dependent, either stopping too early or too late, and fail to account for the expected value of future experiments.
A more principled approach is to treat stopping as part of the sequential decision-making process, naturally leading to a Bayesian optimal stopping formulation that weighs the immediate reward of terminating against the expected value of continuing. While optimal stopping theory is well developed in stochastic processes (e.g., Peskir and Shiryaev 2006), its integration with Bayesian experimental design remains largely unexplored. Existing sequential design methods typically assumes fixed horizons (see Section 2 for a review), leaving open a fundamental gap: no general framework jointly optimizes design and stopping in sequential Bayesian experimental design.
@@
In this work, we close this gap by developing a rigorous framework for optimal stopping in sequential Bayesian experimental design. Our contributions are:
@@
  • We introduce an actor-critic policy gradient algorithm for jointly learning design and stopping policies.
  • We identify a circular dependency in joint optimization that destabilizes training, and resolve it via curriculum learning that gradually transitions from forced continuation to adaptive stopping.
  • Through a linear-Gaussian benchmark and a contaminant source detection problem, we demonstrate that our policy gradient method with curriculum learning achieves stable convergence and outperforms vanilla variants in settings with strong sequential dependencies.
We bridge optimal stopping with Bayesian experimental design, advancing the development of autonomous and resource-efficient experimentation.
2 Related Work
Bayesian experimental design (Chaloner and Verdinelli, 1995; Ryan et al., 2016; Alexanderian, 2021; Rainforth et al., 2024; Huan et al., 2024) provides a principled framework for selecting experiments by maximizing expected information gain (Lindley, 1956). Modern sequential extensions (Foster et al., 2021; Ivanova et al., 2021; Blau et al., 2022; Shen and Huan, 2023; Shen et al., 2025) adapt each experiment to previous data, but almost always assume a fixed horizon: the total number of experiments is set in advance, with no explicit stopping policy.
When early termination is considered, it is usually via simple thresholds, such as halting when posterior variance drops below a cutoff or when budget is exhausted. Similar rules appear in active learning (Zhu et al., 2010; Pullar-Strecker et al., 2024), multi-armed bandits (Audibert and Bubeck, 2010), and online A/B testing (Daskalakis and Kawase, 2017). These methods are convenient but myopic, ignoring the expected long-term value of further experimentation (Figure 1).
Optimal stopping theory is well studied in stochastic processes and dynamic programming (Peskir and Shiryaev, 2006; Shiryaev, 2008; Haggstrom, 1966; Bertsekas, 2012; Puterman, 2014). Reinforcement learning can learn stopping rules but assume a fixed data-acquisition process and do not co-optimize experimental design (Fathan and Delage, 2021; Li and Lee, 2023). Some Bayesian designs include ad hoc stopping; for example, Berry et al. (2002) introduced a “terminator” action in adaptive drug trials using low-dimensional Gaussian approximations and backward induction over small discrete state and action spaces. Their method is specialized and non-information-theoretic, unlike our general MDP formulation with explicit stopping theory and policy gradient training.
Despite progress in sequential design and stopping separately, to our knowledge no prior work provides a general framework for jointly optimizing design and stopping in information-theoretic sequential Bayesian experimental design. Our work closes this gap with a principled MDP formulation, an explicit optimal stopping rule, and a stable learning algorithm.
@@ 3.1 Preliminaries
We consider a finite sequence of $N$ experiments indexed by $k=0,1,\ldots,N-1$. Let $\theta\in\mathbb{R}^{N_{\theta}}$ denote model parameters, $\xi_{k}\in\Xi_{k}\subseteq\mathbb{R}^{N_{\xi}}$ the design for experiment $k$, and $y_{k}\in\mathbb{R}^{N_{y}}$ the corresponding observation. The information history at stage $k$ is $I_{k}=\{\xi_{0},y_{0},\dots,\xi_{k-1},y_{k-1}\}$ with $I_{0}=\emptyset$. The stage-$k$ belief (the prior for experiment $k$) is $p(\theta|I_{k})$. Observations are assumed to follow
@@
where $G_{k}$ is the forward map (which may depend on $I_{k}$), and $\epsilon_{k}$ is an addictive observation noise with known density $p_{\epsilon}$. We assume $\epsilon_{k}$ is conditionally independent across stages given $(\theta,\xi_{k},I_{k})$.
Upon observing $y_{k}$, the belief is updated via Bayes’ rule:
@@
where $p(y_{k}|\theta,\xi_{k},I_{k})$ is the likelihood induced by $(G_{k},p_{\epsilon})$, and $p(y_{k}|\xi_{k},I_{k})=\int p(y_{k}|\theta,\xi_{k},I_{k})p(\theta|I_{k})\,\mathrm{d}\theta$ is the marginal likelihood. The posterior $p(\theta|y_{k},\xi_{k},I_{k})$ then becomes the prior for the next stage, i.e., $p(\theta|I_{k+1})$. This recursive update defines the belief-state dynamics that will underlie our MDP formulation in section 3.2.
@@ 3.2 Markov Decision Process for Optimal Stopping
We formulate optimal stopping for sequential Bayesian experimental design as an MDP (see Figure 2). This extends existing MDP-based sequential design framework (e.g., Shen and Huan, 2023) by introducing an explicit stopping action that enables early termination of the experimental sequence.
@@
States. At stage $k$, the state $s_{k}\in\mathcal{S}$ is represented as $s_{k}=\{s_{k}^{b},s_{k}^{p}\}$. The belief state $s_{k}^{b}$ summarizes the uncertainty about $\theta$, which is determined from $I_{k}$ (for this paper, we can effectively view $s_{k}=I_{k}$). The physical state $s_{k}^{p}$ captures any additional deterministic variables relevant to design (e.g., sensor positions). An additional special terminal state $T$ indicates that the experimental sequence has stopped and no future experiments will be conducted.
@@
Rewards. The reward encodes information gain and experimental cost. Let $r_{k}(s_{k},\xi_{k},y_{k})$ denote the immediate reward at stage $k$, and $r_{T}(s_{k})$ the terminal reward collected when the experimental sequence ends. Experiment $k$ incurs cost $c_{k}(\xi_{k})$. Information gain is measured using the Kullback–Leibler (KL) divergence between distributions. We consider two equivalent reward formulations:
@@
$\pi^{*},$ $\psi^{*}=\operatorname*{arg\,max}_{\pi,\psi}U(\pi,\psi)$ (8)
subject to $\qquad\varphi_{k}(s_{k})\in\{0,1\},\quad\xi_{k}=\mu_{k}(s_{k})\in\Xi_{k},$
$\qquad s_{k+1}=\begin{cases}\mathcal{F}_{k}(s_{k},\xi_{k},y_{k}),&s_{k}\neq T\text{ and }\varphi_{k}(s_{k})=0,\\ T,&\text{otherwise}.\end{cases}$
The objective (expected utility) is the expected total reward:
$U(\pi,\psi)=\mathbb{E}_{y_{0:N-1}\mid\pi,\psi,s_{0}}\left[\sum_{k=0}^{N-1}\left(r_{k}(s_{k},\xi_{k},y_{k})+r_{T}(s_{k})\right)+r_{T}(s_{N})\right].$ (9)
Although the terminal and incremental formulations differ in reward structure, they induce equivalent optimal policies when maximizing expected total reward. This MDP formulation thus unifies design and stopping, providing a principled framework for resource-efficient experimentation.
@@ 4.1 Optimal stopping policy
Before presenting the optimal stopping rule, we introduce the value function (V-function). The V-function at stage $k$ ($k=0,\ldots,N-1$) under design policy $\pi$ and stopping policy $\psi$ is:
@@
which represents the expected cumulative reward from state $s_{k}$ onward when following $\pi$ and $\psi$.
Let us further define $r^{S}_{T}(s_{k})$ to be $r_{T}(s_{k})$ when $s_{k}\neq T$ and $\varphi(s_{k})=1$—i.e., the terminal reward obtained by stopping at state $s_{k}$; and $r^{C}_{k}(s_{k},\mu_{k}(s_{k}),y_{k})$ to be $r_{k}(s_{k},\mu_{k}(s_{k}),y_{k})$ when $s_{k}\neq T$ and $\varphi(s_{k})=0$—i.e., the immediate reward obtained by continuing.
Theorem 1 (Optimal Stopping Policy).
@@
$\varphi_{k}(s_{k})=\bm{1}_{s_{k}\in\mathcal{T}_{k}},\quad k=0,\ldots,N-1,$ (11)
@@
$\mathcal{T}_{k}=\left\{s_{k}\,\Big|\,r^{S}_{T}(s_{k})\geq\mathbb{E}_{y_{k}\mid s_{k},\pi}\left[r^{C}_{k}(s_{k},\mu_{k}(s_{k}),y_{k})+V^{\pi,\psi}_{k+1}\big(\mathcal{F}_{k}(s_{k},\mu_{k}(s_{k}),y_{k})\big)\right]\right\}.$ (12)
We provide a proof in Appendix A.1. Intuitively, the theorem states that the optimal decision at each stage is to compare the terminal reward from stopping with the expected continuation value. Stopping is optimal whenever the former dominates, balancing the value of ending experimentation now against the potential information gain from further experiments.
@@
Thus the Q-function provides a direct way to evaluate whether stopping and taking the terminal reward is better than the value of continuation.
Theorem 2 (Terminal-Incremental Equivalence).
Let $U_{T}(\pi,\psi)$ denote the expected utility under the terminal formulation, and $U_{I}(\pi,\psi)$ the expected utility under the incremental formulation. Then for any policies $(\pi,\psi)$, $U_{T}(\pi,\psi)=U_{I}(\pi,\psi)$.
We provide a proof in Appendix A.2. Hence both reward structures yield the same optimization problem, and one may choose whichever is more convenient for computation.

### v2

@@ Abstract
Sequential Bayesian experimental design typically assumes that the number of experiments is fixed before data collection begins. In practical campaigns, however, experimentation may need to terminate early because additional measurements can provide diminishing information relative to their cost, raising the central decision question: when should one stop? Common threshold-based stopping rules are easy to implement but myopic, because they compare the current state with a fixed criterion without accounting for the expected value of future experiments. This work develops a Bayesian optimal stopping framework for sequential experimental design by formulating stopping and design as coupled decisions in a Markov decision process. We prove that, for any design policy, the optimal stopping rule terminates exactly when the immediate terminal reward exceeds the expected continuation value. We then derive a policy gradient method for learning value-based stopping and design policies. Naïve joint training can create a circular dependency that traps learning in early-stopping local optima. We address this difficulty with a curriculum learning strategy that gradually transitions from forced continuation to adaptive stopping during training. Numerical studies on a linear-Gaussian benchmark, a one-dimensional nonlinear test problem, and a contaminant source detection problem show that the proposed approach learns stable design–stopping policies and improves resource-aware performance, with the largest gains in settings with strong sequential dependence.
Keywords Bayesian experimental design $\cdot$ optimal stopping $\cdot$ Markov decision process $\cdot$ policy gradient $\cdot$ curriculum learning $\cdot$ sequential design
@@ 1 Introduction
Sequential experimentation plays a central role in science and engineering, from clinical trials (Murphy, 2003) and materials discovery (Lookman et al., 2019) to environmental monitoring (Krause et al., 2008). In these settings, practitioners adaptively choose each new experiment using the outcomes of previous experiments. A fundamental but often overlooked question is therefore inseparable from design itself: when should experimentation stop?
@@
Bayesian experimental design (Chaloner and Verdinelli, 1995; Ryan et al., 2016; Alexanderian, 2021; Rainforth et al., 2024; Huan et al., 2024) provides a principled framework for selecting experiments by maximizing expected information gain (Lindley, 1956). Modern sequential extensions (Foster et al., 2021; Ivanova et al., 2021; Blau et al., 2022; Shen and Huan, 2023; Shen et al., 2025) adapt each experiment to previously collected data, but they usually assume a fixed experimental horizon: the total number of experiments is specified in advance and no explicit stopping policy is optimized. This assumption is restrictive in practice, where campaigns often face uncertain budgets, limited resources, or diminishing returns that make early termination desirable. Existing early-termination strategies are commonly based on thresholds that stop once a predefined statistic, such as posterior variance or remaining budget, crosses a cutoff. Related threshold rules also appear in active learning (Zhu et al., 2010; Pullar-Strecker et al., 2024), multi-armed bandits (Audibert and Bubeck, 2010), and online A/B testing (Daskalakis and Kawase, 2017). Although convenient, these rules are inherently myopic: they depend only on the current state and ignore the expected long-term trade-off between the value of additional information and the cost of further experiments. Consequently, a threshold may terminate too early, missing valuable information, or too late, wasting resources (see Figure 1).
@@
Figure 1: Illustrative accumulated reward trajectories over six experiments, generated from the linear-Gaussian benchmark, showing the limitations of threshold-based stopping. Here, a fixed threshold of 1.8 (dashed line) yields very different outcomes across trajectories: the orange path is stopped prematurely after one experiment, missing the higher rewards achievable at stages 2–3; the blue path never crosses the threshold, leading to unnecessary continuation past its peak; the green path happens to cross the threshold near its maximum, but only by chance. Threshold rules are thus highly outcome-dependent, either stopping too early or too late, and fail to account for the expected value of future experiments.
A more principled approach is to make stopping part of the sequential decision problem. This viewpoint leads naturally to a Bayesian optimal stopping formulation in which the immediate reward of terminating is compared with the expected value of continuing. Optimal stopping theory is well developed in stochastic processes and dynamic programming (Peskir and Shiryaev, 2006; Shiryaev, 2008; Haggstrom, 1966; Bertsekas, 2012; Puterman, 2014). Existing reinforcement learning formulations of stopping typically use a binary action space {continue, stop}, learning a stopping policy without jointly optimizing the experimental design or control action selected upon continuation (Fathan and Delage, 2021; Li and Lee, 2023). Some Bayesian designs include ad hoc termination mechanisms; for example, Berry et al. (2002) introduced a “terminator” action in adaptive drug trials using low-dimensional Gaussian approximations and backward induction over small discrete state and action spaces. That approach is valuable but specialized and not information-theoretic. Related ideas also appear in Bayesian optimization (BO), where the Knowledge Gradient (KG) (Frazier et al., 2008; Ryzhov et al., 2012) continues sampling only when the expected one-step improvement exceeds evaluation cost (Garnett, 2023; Xie et al., 2025). While effective for BO objectives, KG remains a one-step criterion tied to optimization rather than a general information-theoretic formulation for sequential Bayesian experimental design with optimal stopping.
Despite progress in sequential design and optimal stopping as separate topics, a general framework is still missing for jointly optimizing design and stopping in information-theoretic sequential Bayesian experimental design. Simply adding a termination action to an existing sequential design method is not enough, because stopping changes the decision problem by introducing an endogenous horizon, terminal rewards, and a state-dependent stopping boundary. These structural changes create both theoretical and computational challenges. Amortized methods such as Deep Adaptive Design (DAD) (Foster et al., 2021) rely on pathwise gradient estimators for fixed-horizon trajectories with smooth dependence on policy parameters, making discrete termination decisions difficult to incorporate. Conversely, generic reinforcement learning approaches (Blau et al., 2022) can represent stopping in discrete settings, but in continuous design spaces they typically rely on score-function estimators and model stopping as a stochastic action rather than a deterministic value-based decision. Appendix A.6 provides an empirical illustration of this limitation using a generic hybrid-action reinforcement learning baseline. More broadly, naïve joint optimization creates a circular dependency: the design policy determines future information, the continuation value determines the stopping boundary, and the stopping boundary determines which trajectories are observed during training. This feedback can cause premature termination and poor sample efficiency.
@@
In this work, we address these challenges by developing a rigorous framework for optimal stopping in sequential Bayesian experimental design. Our contributions are:
@@
  • We introduce a policy gradient algorithm with value-function approximation for learning continuous design policies with value-based stopping.
  • We identify a circular dependency that can trap training in early-stopping local optima, and mitigate it using curriculum learning that gradually transitions from forced continuation to adaptive stopping.
  • Through a linear-Gaussian benchmark, a one-dimensional nonlinear test problem, and a contaminant source detection problem, we demonstrate stable, resource-aware design–stopping policies, with the largest curriculum gains in settings with strong sequential dependence.
Together, these contributions connect optimal stopping with Bayesian experimental design and provide a foundation for autonomous, resource-aware experimentation.
This paper is structured as follows. Section˜2 formulates sequential Bayesian experimental design with stopping as a Markov decision process. Section˜3 derives the optimal stopping rule, develops the policy gradient algorithm, and introduces curriculum learning for stabilizing training under circular design–stopping dependencies. Section˜4 validates the approach on a linear-Gaussian benchmark, a one-dimensional nonlinear test problem, and a contaminant source detection problem. Section˜5 concludes.
@@ 2.1 Preliminaries
We consider a finite sequence of $N$ experiments indexed by $k=0,1,\ldots,N-1$. Let $\theta\in\mathbb{R}^{N_{\theta}}$ denote model parameters, $\xi_{k}\in\Xi_{k}\subseteq\mathbb{R}^{N_{\xi}}$ the design for experiment $k$, and $y_{k}\in\mathbb{R}^{N_{y}}$ the corresponding observation. The information history at stage $k$ is $I_{k}=\{\xi_{0},y_{0},\dots,\xi_{k-1},y_{k-1}\}$ with $I_{0}=\emptyset$. The stage-$k$ belief (i.e., the prior for experiment $k$) is $p(\theta\mid I_{k})$. Observations are assumed to follow
@@
where $G_{k}$ is the forward map (which may depend on $I_{k}$), and $\epsilon_{k}$ is an additive observation noise with known density $p_{\epsilon}$. We assume $\epsilon_{k}$ is conditionally independent across stages given $(\theta,\xi_{k},I_{k})$. Upon observing $y_{k}$, the belief is updated via Bayes’ rule:
@@
where $p(y_{k}|\theta,\xi_{k},I_{k})$ is the likelihood induced by $(G_{k},p_{\epsilon})$, and $p(y_{k}|\xi_{k},I_{k})=\int p(y_{k}|\theta,\xi_{k},I_{k})p(\theta|I_{k})\,\mathrm{d}\theta$ is the marginal likelihood. The posterior $p(\theta|y_{k},\xi_{k},I_{k})$ then becomes the prior for the next stage, i.e., $p(\theta|I_{k+1})$. This recursive update defines the belief-state dynamics that will underlie our MDP formulation in Section˜2.2.
@@ 2.2 Markov Decision Process for Optimal Stopping
We formulate optimal stopping for sequential Bayesian experimental design as an MDP (see Figure 2). This extends existing MDP-based sequential design frameworks (e.g., Shen and Huan, 2023) by introducing an explicit stopping action that enables early termination of the experimental sequence.
@@
States. At stage $k$, the state $s_{k}\in\mathcal{S}$ is represented as $s_{k}=\{s_{k}^{b},s_{k}^{p}\}$. The belief state $s_{k}^{b}$ summarizes the uncertainty about $\theta$ and is fully determined by $I_{k}$; for this paper, we can equivalently view $s_{k}=I_{k}$, yielding a Markov state representation. The physical state $s_{k}^{p}$ captures any additional deterministic variables relevant to design (e.g., sensor positions). An additional special terminal state $T$ indicates that the experimental sequence has stopped and no future experiments will be conducted.
@@
This transition structure induces an endogenous (state-dependent) horizon determined by the stopping policy.
Rewards. The reward encodes information gain and experimental cost. Let $r_{k}(s_{k},\xi_{k},y_{k})$ denote the immediate reward at stage $k$, and $r_{T}(s_{k})$ the terminal reward collected when the experimental sequence ends. Experiment $k$ incurs a (negative) cost $c_{k}(\xi_{k})$. Information gain is measured using the Kullback–Leibler (KL) divergence between distributions. We consider two equivalent reward formulations:
@@
$\pi^{*},$ $\hskip 3.00003pt\psi^{*}\in\operatorname*{arg\,max}_{\pi,\psi}U(\pi,\psi)$ (8)
subject to
$\varphi_{k}(s_{k})\in\{0,1\},\quad\xi_{k}=\mu_{k}(s_{k})\in\Xi_{k},$
$s_{k+1}=\begin{cases}\mathcal{F}_{k}(s_{k},\xi_{k},y_{k}),&s_{k}\neq T\text{ and }\varphi_{k}(s_{k})=0,\\ T,&\text{otherwise}.\end{cases}$
where the objective (expected utility) is the expected total reward:
$U(\pi,\psi)=\mathbb{E}_{y_{0:N-1}\mid\pi,\psi,s_{0}}\Bigg[\sum_{k=0}^{N-1}\left(r_{k}(s_{k},\xi_{k},y_{k})+r_{T}(s_{k})\right)+r_{T}(s_{N})\Bigg].$ (9)
Although the terminal and incremental formulations differ in reward structure, they induce equivalent optimal policies when maximizing expected total reward. This MDP formulation unifies design and stopping as joint decision variables, providing a principled framework for resource-efficient experimentation.
@@ 3 Solution Method
This section develops the computational framework for learning sequential designs with optimal stopping. We first characterize the value-based stopping rule induced by any fixed design policy. We then use this structure to reduce joint design–stopping optimization to a policy gradient update over the continuous design policy, with a critic network estimating the continuation value. Finally, we introduce a curriculum strategy that stabilizes training by delaying the influence of learned stopping decisions until the design and value networks have observed sufficiently informative trajectories.
@@ 3.1 Optimal stopping policy
We begin by defining the value function used to compare stopping and continuation. The V-function at stage $k$ ($k=0,\ldots,N-1$) under design policy $\pi$ and stopping policy $\psi$ is:
@@
which represents the expected cumulative reward from state $s_{k}$ onward when following $\pi$ and $\psi$. Equivalently, the value function satisfies the recursion $V^{\pi,\psi}_{k}(s_{k})=\mathbb{E}_{y_{k}\mid\pi,\psi,s_{k}}\big[r_{k}(s_{k},\mu_{k}(s_{k}),y_{k})+r_{T}(s_{k})+V^{\pi,\psi}_{k+1}(s_{k+1})\big]$ with $V^{\pi,\psi}_{N}(s_{N})=r_{T}(s_{N})$.
Define $r_{T}^{S}(s_{k})$ as the terminal reward obtained by stopping at $s_{k}$ (i.e., $s_{k}\neq T$ and $\varphi_{k}(s_{k})=1$), and $r_{k}^{C}(s_{k},\mu_{k}(s_{k}),y_{k})$ as the immediate reward obtained by continuing (i.e., $s_{k}\neq T$ and $\varphi_{k}(s_{k})=0$).
Theorem 3.1 (Optimal Stopping Policy).
@@
$\varphi_{k}(s_{k})=\mathbf{1}_{s_{k}\in\mathcal{T}_{k}},\quad k=0,\ldots,N-1,$ (11)
@@
$\mathcal{T}_{k}=\Big\{s_{k}\ \Big|\ r^{S}_{T}(s_{k})\geq$ $\mathbb{E}_{y_{k}\mid s_{k},\pi}\Big[r^{C}_{k}(s_{k},\mu_{k}(s_{k}),y_{k})+V^{\pi,\psi}_{k+1}\big(\mathcal{F}_{k}(s_{k},\mu_{k}(s_{k}),y_{k})\big)\Big]\Big\}.$ (12)
We provide a proof in Appendix A.1. The theorem states that the optimal decision at each stage is a value comparison: stop if the terminal reward is no smaller than the expected value of conducting one more experiment and then acting optimally. This rule balances the value of ending experimentation immediately against the potential information gain from future experiments.
@@
Thus the Q-function provides a direct criterion for deciding whether the terminal reward is preferable to the value of continuation.
Theorem 3.2 (Terminal-Incremental Equivalence With Stopping).
Let $U_{T}(\pi,\psi)$ denote the expected utility under the terminal formulation in (4) and (5), and $U_{I}(\pi,\psi)$ the expected utility under the incremental formulation in (6) and (7). Then for any policies $(\pi,\psi)$, $U_{T}(\pi,\psi)=U_{I}(\pi,\psi)$.
We provide a proof in Appendix A.2. While terminal–incremental equivalence is known for fixed-horizon sequential design (Foster et al., 2021; Shen and Huan, 2023), extending this result to settings with policy-dependent stopping requires explicitly proving consistency of optimal stopping across both formulations. This extension is important computationally because it ensures that both reward formulations yield equivalent optimal policies while allowing implementation choices that trade off reward sparsity against per-stage posterior computations. We return to this tradeoff in Section 3.2.